OpenStudy (lilthesis101):

A forest Ranger spots a fire from a 28 foot tower. The angle of depression from the tower to the fire is 11°. To the nearest foot, how far is the fire from the base of the tower? Show the steps you used to find the solution.

OpenStudy (lilthesis101):

@undeadknight26

OpenStudy (lilthesis101):

@ganeshie8

OpenStudy (oleg3321):

Now, since the angle of depression is 11°, we can know the angle inside the triangle is 79°, since 11+79=90 (Forms right angle). Now, since a right angle is 90°, and since all angles in a triangle sums up to 180° (Triangle Angle Sum Theorem), the angle of elevation is 11°. Now we can use the Law of Sines to find the answer ("x" is the side we are looking for). x/sin(90) = 28/ sin(11) x = (28/sin(11))sin(90) And, using a calculator: x= 147 (rounded to the nearest foot) So the distance between the tower and the fire is 147 feet. Hope this helps!

jimthompson5910 (jim_thompson5910):

 T = top of tower B = base of tower F = location of fire use the trig function tangent to solve for x tan(11 degrees) = 28/x